OpenStudy (anonymous):

find the area bounded by the curve {(x,y): |x-1| <= y <= sqrt (5- x^2)}

OpenStudy (anonymous):

are these the parameters \[\left| x-1\right|\le y \le \sqrt{5}-x^{2}\]?

OpenStudy (anonymous):

yes

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

any ans pls???

OpenStudy (anonymous):

Maybe it would help to draw the area you have to find (ignore the bottom half of the circle):